ExERCISE 3 In 1980, New York City had a more-or-less free market of taxi services. Any firm could provide a taxi service, as long as the drivers and cars satisfied certain safety standards Suppose the constant marginal cost per trip of a taxi ride is $5, and the average taxi has a capacity of 20 trips per day. Let the demand function for taxi rides be given by D(p) 1,200-20p, where demand is measured in rides per day, and price is measured in dollars. Assume that the industry is perfectly competitive. (1) What is the competitive equilibrium price per ride? What is the equilibrium number of rides per day? How many taxicabs there will be in equilibrium? (2) In 1980 the city council created a licensing board and issued a license to each cab. Then, they stated that it would continue to adjust taxicab fares so that the demand for rides equals the supply, but no new licenses will be issued in the future. In 1985 the demand for rides is D(p) 1.220-20p. What was the equilibrium price of a ride in 1985? (3) What was the profit per ride in 1985? And per taxicab license per day? If the taxi operated every day, what was the profit per license per year? (4) If the interest rate was %10 a year and costs, demand and the number of licenses remain constant forever, what would be the market price of a taxicab license?
Answer #1

Hello,

Please find the answer below:

1 a At perfectly competitive equilibrium price = MC. Here MC = $5. So perfectly competitive equi price = $5

b As P= 5 , D = 1200 -20p = 1200 - 20*5 = 1200-100 = 1100 equilibrium number of rides.

c One taxi can make 20 trips. So for a demand of 1100 rides , we need 1100/20 = 55 taxis.

2 Demand equals supply. Thus, the supply of rides is 1100.

and new demand D(p)= 1220-20p

1100=1220-20p

p= $6

3 a MC per ride = $5 and fare is $6

Profit per ride = 6-5 = $1

b Yearly profit = $1 * 20 rides * 365 days = 7300

4 Interest rate = 10%

when everything remains constant and profit = 7300

let market price = x

0.1 x = 7300

= 7300/0.1

= $73000

Hope this helps you. IT will be great if you can share your feedback. Thank you!!